OpenStudy (anonymous):

@campbell_st @aum @deshawn1 Can you help me solve this problem please. Ian has 10 fewer dimes than quarters. He has $5.30. Find the numbers of quarters and dimes Ian has.

OpenStudy (anonymous):

the two equations are D + 10 = Q i.e 1) Q-D =10 2)10D+25Q=530 multiply equation 1 with 10 by using elimination method we get 10D - 10Q =100 10D + 25Q = 530 _____________________ 35Q= 630 therefore, Q= 18 then substitute Q in equation 1 to get D+10=Q, ie, D=18-10=8.